the volume occupied by 150 grams of chlorine at -12.5 Celsius and 98.5 kPa


1
Expert's answer
2022-03-01T14:44:02-0500

Number of moles= 150g/70.906 = 2.12 moles.

T= -12.5+273=260.5k

R= 0.0821

P= 0.972atm so Pv= nRT

0.972v=2.12x0.0821×26.5

0.972v= 4.612378

v= 4.745244856